# arXiv:2603.01198v1[eess.SY] 01 Mar 2026

### v1

@@ 1 Introduction
Global data center electricity consumption reached 415 TWh in 2024, approximately 1.5% of worldwide electricity demand, and is projected to exceed 945 TWh by 2030 (International Energy Agency, 2025). In the United States alone, data centers consumed 176 TWh in 2023 (4.4% of national electricity), with projections ranging from 325 to 580 TWh by 2028 (Shehabi et al., 2024). While advances in server efficiency and virtualization have partially offset demand growth (Masanet et al., 2020), the proliferation of artificial intelligence workloads and the push toward exascale computing are placing unprecedented strain on cooling infrastructure. Cooling systems account for 30–40% of total data center electricity consumption (Zhang et al., 2021; Ebrahimi et al., 2014), making them the single largest controllable load in most facilities. Yet the global average Power Usage Effectiveness (PUE) has stagnated between 1.55 and 1.59 since 2020 (Uptime Institute, 2024), suggesting that incremental improvements have plateaued and that fundamentally new optimization approaches are needed.
High-performance computing (HPC) facilities intensify this challenge. Exascale supercomputers such as Frontier at Oak Ridge National Laboratory operate at power levels of 8–30 MW, employ 100% direct liquid cooling with variable-speed pumps circulating thousands of liters per minute, and experience rapid thermal transients driven by workload dynamics (Sun et al., 2024; Karimi et al., 2024). These systems differ structurally from enterprise data centers: cooling loops have response times measured in minutes, thermal loads fluctuate on the scale of seconds, and equipment constraints such as pump ramp-rate limits are operationally binding. Despite the scale of energy involved, systematic cooling system optimization for HPC facilities remains largely unexplored. Existing work on Frontier has focused on operational data collection (Sun et al., 2024), power management (Karimi et al., 2024), thermal stability (Grant et al., 2026), and waste heat recovery (Wang et al., 2024). A data-driven machine learning framework for Frontier was developed to quantify cooling inefficiency using gradient-boosting surrogate models, identifying approximately 85 MWh of annual excess cooling energy (Jadhav and Liu, 2026). However, no study has coupled a physics-based digital twin of an exascale cooling system with systematic optimization, nor has any work formally quantified how much of the theoretical energy savings survive when practical operational constraints are imposed.
This paper addresses that question through three co-equal contributions: (i) a validated Modelica-based digital twin of Frontier’s liquid cooling system, (ii) a layered optimization framework using classical gradient-based methods that progressively reveals the sources and magnitudes of energy savings, and (iii) a formally defined implementability gap metric that quantifies the energy savings lost when theoretical optima must accommodate practical operational constraints. The remainder of this section reviews the relevant literature across four domains and positions the present work within identified research gaps.
@@ 1.1 Data center cooling system optimization: approaches and limitations
Research on data center cooling system optimization has accelerated considerably since a 40% reduction in Google’s cooling energy was reported using neural network recommendations (Evans and Gao, 2016). Although that claim has been critiqued for its lack of peer review and reproducibility (Lazic et al., 2018), it catalyzed substantial academic and industrial interest in intelligent cooling control.
Reinforcement learning (RL) approaches treat the cooling system as a Markov decision process and learn control policies from interaction with the environment or a simulator. Deep RL has been demonstrated for data center cooling with promising results on simulated workloads (Li et al., 2020), and a multi-agent RL framework integrated with a digital twin has shown potential for coordinated cooling control (Sarkar et al., 2024). More recently, 14–21% energy savings were reported from offline RL deployed in a production data center, representing one of the few validated real-world implementations (Zhan et al., 2025). RL methods offer the advantage of learning directly from data without requiring an explicit system model, but they introduce challenges around sample efficiency, safety during exploration, and interpretability of the learned policy, all of which are particularly acute in safety-critical HPC environments. More broadly, data-driven surrogates (including gradient boosting models) have been applied to diagnose cooling inefficiency at HPC facilities (Jadhav and Liu, 2026), but these approaches identify waste rather than prescribing constrained optimal control trajectories.
Model predictive control (MPC) formulates cooling system optimization as a receding-horizon constrained optimization problem, typically requiring a dynamical model of the thermal system. MPC has been applied to Google’s data centers, achieving a rigorously documented 17.9% cost reduction (Lazic et al., 2018). A foundational framework for stochastic MPC in building climate control was established for commercial buildings (Oldewurtel et al., 2012), and MPC has been demonstrated specifically for chiller plant operation (Ma et al., 2012). MPC naturally handles constraints and can incorporate weather forecasts, but its performance degrades substantially with model uncertainty. Simple rule-based control has been shown to outperform MPC when model uncertainty exceeds approximately 67% (Maasoumy et al., 2014), highlighting a practical limitation that is often overlooked.
Classical gradient-based optimization methods such as Sequential Quadratic Programming (SQP) solve constrained nonlinear programs using gradient information and are the workhorse of engineering design optimization (Kraft, 1988). These methods are deterministic, computationally transparent, and handle equality and inequality constraints through well-understood Lagrangian and active-set strategies. For cooling systems governed by smooth physics (e.g., the cubic affinity law relating pump power to flow rate), the optimization landscape is well suited to gradient methods. Despite these advantages, classical gradient-based optimization is notably absent from the data center and HPC cooling literature. The field has overwhelmingly favored RL and MPC, leaving a significant methodological gap.
@@
Table 1 summarizes the positioning of representative works across these three families, together with the present study. No existing paper simultaneously addresses physics-based digital twin modeling, systematic optimization, HPC liquid cooling, explicit practical constraints, and a formal implementability metric.
Table 1: Comparison of representative cooling system optimization studies. Abbreviations: DT = digital twin; HPC = high-performance computing; DC = data center.
Study Method Application DT Constraints Gap metric
Evans and Gao (2016) Neural net Google DC No No No
Lazic et al. (2018) MPC Google DC No Partial No
Li et al. (2020) Deep RL Enterprise DC No No No
Sarkar et al. (2024) Multi-agent RL Enterprise DC Yes No No
Zhan et al. (2025) Offline RL Production DC No Partial No
Zhu and Lin (2024) Real-time opt. Enterprise DC Yes Partial No
Jadhav and Liu (2026) ML surrogate Frontier (HPC) No No No
Brewer et al. (2024) — (V&V only) Frontier (HPC) Modelica No
Fu et al. (2019a) — (Sim. only) Air-cooled DC Modelica No
This study SLSQP (layered) Frontier (HPC) Modelica Ramp-rate Yes
@@ 1.2 Digital twins for thermal system modeling and optimization
The digital twin concept, originally articulated for product lifecycle management (Grieves and Vickers, 2017) and subsequently formalized for industrial applications (Tao et al., 2019), has found growing adoption in building and data center energy systems. The IISE community has recognized digital twins as a transformative methodology for design and manufacturing (Knapp et al., 2020; Tan and Matta, 2024), with active research spanning synchronization, fidelity, and optimization. Digital twins for thermal systems generally follow one of three paradigms: physics-based models that encode conservation laws and constitutive relations, data-driven models that learn input-output mappings from operational data, and hybrid approaches that combine both (Afram and Janabi-Sharifi, 2014). For optimization applications where control strategies must be explored beyond the range of historical operating data, physics-based models offer superior generalization and physical interpretability.
@@
The Modelica language and its associated Buildings Library (Wetter et al., 2014) have emerged as the dominant open-source platform for equation-based modeling of HVAC and cooling systems. This library provides validated component models for heat exchangers, pumps, valves, pipes, and control systems that can be composed into arbitrarily complex system topologies. The first comprehensive Modelica-based data center cooling models were developed for a conventional air-cooled facility with computer room air handlers (Fu et al., 2019a, b). Modelica modeling has been extended to district cooling systems (Hinkelman et al., 2022), and open-source Modelica models for chiller plants with water-side economizers have reported energy savings of 8.6–36.8% across operating scenarios (Fan et al., 2021). More recently, Modelica has been used for comparative co-design of data center cooling controls (Grahovac et al., 2023), and the Buildings Library has been coupled with EnergyPlus through the Spawn framework (Wetter et al., 2024).
@@
For HPC cooling specifically, the Exascale Digital Twin (ExaDigiT) project at Oak Ridge National Laboratory developed a digital twin framework for liquid-cooled supercomputers using the TRANSFORM library within the Modelica environment (Brewer et al., 2024; Kumar et al., 2024; Greenwood et al., 2024). ExaDigiT models the thermo-fluid dynamics of Frontier’s cooling loops and has been validated against operational sensor data. However, the ExaDigiT effort focused on verification and validation of the digital twin itself; it did not proceed to use the twin for systematic cooling optimization. This pattern recurs across the literature: several groups have built physics-based cooling models (Fu et al., 2019a; Hinkelman et al., 2022) or digital twin frameworks (Athavale et al., 2024; Sarkar et al., 2024), but the step from validated model to actionable optimization is frequently omitted or treated superficially.
The present work constructs an independent Modelica-based digital twin of Frontier’s cooling system, distinct from ExaDigiT in both modeling approach and purpose. ExaDigiT employs the TRANSFORM library and targets broad digital twin capabilities including real-time monitoring and what-if analysis; the model developed here uses the Modelica Buildings Library (Wetter et al., 2014) and is purpose-built for steady-state optimization. The model is calibrated against 47,186 validated operational records from a published operational dataset (Sun et al., 2024) and validated to ASHRAE Guideline 14 (ASHRAE, 2014) criteria, achieving coefficient of variation of the root mean square error (CV-RMSE) values of 1.96–2.67% and normalized mean bias error (NMBE) within $\pm$10% across all monitored variables (Section 2.3). This level of validation provides the confidence necessary to trust optimization results derived from the model.
@@ 1.3 The implementability gap: from theoretical optima to practical control
A persistent challenge in cooling system optimization is the gap between what is theoretically optimal and what can be practically implemented. This gap arises from multiple sources: actuator limitations (pumps and valves cannot change state instantaneously), thermal inertia (large fluid volumes resist rapid temperature changes), safety margins (component temperatures must remain within rated bounds during transients), and operational protocols (facility operators require gradual, predictable system behavior). Despite the practical importance of these constraints, the literature lacks a formal framework for quantifying their effect on achievable energy savings.
The closest conceptual precedents come from three distinct and largely disconnected research areas. In building energy research, the performance gap between predicted and measured energy consumption has been extensively documented, with discrepancies of 20–550% reported across building types (de Wilde, 2014; Van Dronkelaar et al., 2016). However, this performance gap refers to the mismatch between design-stage predictions and as-built operation, not between optimal and implementable control strategies. In robotics and reinforcement learning, the simulation-to-real gap(also called the reality gap) describes the degradation in controller performance when policies trained in simulation are transferred to physical systems (Salvato et al., 2021; Zhao et al., 2020). This has been addressed specifically for building HVAC RL, but without defining a quantitative metric (Chen et al., 2020). In energy storage optimization, imposing ramp-rate constraints of 10% of maximum capacity has been shown to reduce arbitrage profit by 10–35% (Hashmi et al., 2024), with constrained solutions retaining 65–90% of unconstrained profit. This represents the closest quantitative analog to our work, but it addresses battery storage rather than cooling systems.
None of these research areas formally defines a metric for the gap between unconstrained optimal control and practically implementable control in cooling systems. We address this by introducing the implementability gap, defined as the fraction of theoretical optimal energy savings that cannot be captured due to practical operational constraints, and its complement, the recovery ratio, which quantifies the fraction of theoretical savings that remain achievable (Section 2.4.4). This framework is operationalized through a layered optimization approach: Strategy A (flow-only optimization) establishes a conservative baseline by optimizing only pump speeds; Strategy B (unconstrained co-optimization) simultaneously optimizes pump speeds and cooling tower fan operation without operational constraints, revealing the theoretical maximum savings; and Strategy C (ramp-constrained co-optimization) re-introduces ramp-rate limits on all actuators, yielding the practically implementable savings. The difference between Strategies B and C directly measures the implementability gap, while the ratio of Strategy C savings to Strategy B savings gives the recovery ratio.
This layered design serves a dual purpose. Methodologically, it isolates the contribution of each optimization degree of freedom and each constraint, providing insight that a single monolithic optimization would obscure. Practically, it gives facility operators a transparent menu of strategies with clearly quantified trade-offs between theoretical savings and implementation complexity.
@@ 1.4 Contributions and paper organization
    A validated Modelica-based digital twin of the Frontier supercomputer’s three-loop liquid cooling system, independently developed using the Modelica Buildings Library, calibrated against one year of operational data (47,186 records), and validated to ASHRAE Guideline 14 criteria with CV-RMSE of 1.96–2.67%.
@@
    A layered optimization framework comprising three strategies of increasing complexity (flow-only, unconstrained co-optimization, and ramp-constrained co-optimization) implemented using SLSQP, representing the first application of classical gradient-based optimization to HPC cooling systems.
@@
    The implementability gap metric and recovery ratio, formal measures of the energy savings lost (and retained) when practical ramp-rate constraints are imposed on theoretically optimal control strategies. To the authors’ knowledge, this is the first such metric defined for cooling system optimization.
@@
    Operational insights for exascale cooling, including the finding that flow-only optimization achieves 20.4% pumping energy savings, unconstrained co-optimization reaches 30.1%, and ramp-constrained co-optimization retains 27.8% (corresponding to a recovery ratio of 92.4%), demonstrating that the majority of theoretical savings are practically recoverable.
The remainder of this paper is organized as follows. Section 2 presents the digital twin model, including the system architecture, governing equations, calibration procedure, and validation protocol. Section 3 presents the validation results, optimization findings, and energy savings across strategies. Section 4 discusses system-level insights, practical implementability, robustness, limitations, and directions for future work. Section 5 concludes with principal findings.
@@ 2.2.1 Heat exchanger model
where $f_{i}$ is the fraction of total flow through subloop $i$. This expression follows directly from the first law of thermodynamics applied to the coolant stream passing through the CDU cluster, under the assumption that all waste heat $Q_{i}$ is absorbed by the HTW fluid. The flow fractions are determined by the parallel hydraulic network topology and remain approximately constant at $f_{1}=0.244$, $f_{2}=0.258$, and $f_{3}=0.498$ across the full operating range, as derived from the operational dataset and confirmed by the Modelica hydraulic model (Section 3).
@@
To enable closed-form optimization (Section 2.4), the heat exchanger thermal resistance is characterized by an aggregate overall heat transfer coefficient-area product $UA$. Under the assumption that the HTW-side convective resistance dominates the overall thermal resistance (i.e., wall conduction and CTW-side resistance are comparatively small), the Dittus-Boelter correlation for turbulent internal flow yields $h\propto\dot{m}^{0.8}$ (Incropera et al., 2007), and consequently:
@@ 3.1 Digital twin validation
Table 2 reports the per-subloop prediction errors. All three subloops satisfy both the ASHRAE Guideline 14 hourly thresholds (CV-RMSE $\leq 30\%$, $|\mathrm{NMBE}|\leq 10\%$) and the stricter targets adopted in this study (CV-RMSE $\leq 5\%$, $|\mathrm{NMBE}|\leq 5\%$). Subloop 3 exhibits the largest CV-RMSE (2.67%) due to its higher mean temperature differential ($\overline{\Delta T}_{3}=13.8\,^{\circ}$C vs. $\overline{\Delta T}_{1,2}\approx 8\,^{\circ}$C), which amplifies sensitivity to the assumed $c_{p}$. The positive NMBE across all subloops indicates a consistent over-prediction bias of 0.5–0.9 C, attributable to the 6% difference between the adopted $c_{p}$ and the dataset’s internal conversion constant.

### v2

# Digital Twin-Based Cooling System Optimization for Data Center

@@ 1 Introduction
Global data center electricity consumption reached 415 TWh in 2024, approximately 1.5% of worldwide electricity demand, and is projected to exceed 945 TWh by 2030 (International Energy Agency, 2025). In the United States alone, data centers consumed 176 TWh in 2023 (4.4% of national electricity), with projections ranging from 325 to 580 TWh by 2028 (Shehabi et al., 2024). While advances in server efficiency and virtualization have partially offset demand growth (Masanet et al., 2020), the proliferation of artificial intelligence workloads and the push toward exascale computing are placing unprecedented strain on cooling infrastructure. Cooling systems account for 30–40% of total data center electricity consumption (Zhang et al., 2021; Ebrahimi et al., 2014), making them the single largest controllable load in most facilities. Yet the global average Power Usage Effectiveness (PUE) has stagnated between 1.55 and 1.59 since 2020 (Uptime Institute, 2024), suggesting that fundamentally new optimization approaches are needed.
High-performance computing (HPC) facilities intensify this challenge. Exascale supercomputers such as Frontier at Oak Ridge National Laboratory operate at power levels of 8–30 MW, employ 100% direct liquid cooling with variable-speed pumps, and experience rapid thermal transients driven by workload dynamics (Sun et al., 2024; Karimi et al., 2024). These systems differ structurally from enterprise data centers: cooling loops have response times measured in minutes, and equipment constraints such as pump ramp-rate limits are operationally binding. Despite the scale of energy involved, systematic cooling system optimization for HPC facilities remains largely unexplored. Existing work on Frontier has focused on operational data collection (Sun et al., 2024), power management (Karimi et al., 2024), thermal stability (Grant et al., 2026), waste heat recovery (Wang et al., 2024), and data-driven cooling inefficiency diagnosis (Jadhav and Liu, 2026). However, no study has coupled a physics-based digital twin of an exascale cooling system with systematic optimization, nor has any work formally quantified how much of the theoretical energy savings survive when practical operational constraints are imposed.
This paper addresses that gap through three contributions: (i) a validated Modelica-based digital twin of Frontier’s liquid cooling system, (ii) a layered optimization framework that progressively reveals the sources and magnitudes of energy savings, and (iii) a formally defined implementability gap metric that quantifies the energy savings lost when theoretical optima must accommodate practical constraints.
@@ 1.1 Data center cooling system optimization: approaches and limitations
Research on data center cooling system optimization has accelerated considerably since a 40% reduction in Google’s cooling energy was reported using neural network recommendations (Evans and Gao, 2016), though that claim was subsequently critiqued for its lack of peer review and reproducibility (Lazic et al., 2018). The resulting body of work spans three broad methodological families. Reinforcement learning (RL) approaches treat the cooling system as a Markov decision process: deep RL has shown promising results on simulated workloads (Li et al., 2020), multi-agent RL integrated with a digital twin has demonstrated coordinated cooling control (Sarkar et al., 2024), and 14–21% energy savings have been reported from offline RL deployed in a production data center (Zhan et al., 2025). However, RL methods face challenges around sample efficiency, safety during exploration, and policy interpretability that are particularly acute in safety-critical HPC environments. Model predictive control (MPC) formulates cooling optimization as a receding-horizon constrained problem; a rigorously documented 17.9% cost reduction was achieved at Google’s data centers (Lazic et al., 2018), and MPC has been applied to building climate control (Oldewurtel et al., 2012) and chiller plant operation (Ma et al., 2012). MPC naturally handles constraints but degrades substantially with model uncertainty, with rule-based control shown to outperform MPC when uncertainty exceeds approximately 67% (Maasoumy et al., 2014). Classical gradient-based optimization, the workhorse of engineering design (Kraft, 1988), is deterministic, computationally transparent, and well suited to cooling systems governed by smooth physics, yet it is notably absent from the data center cooling literature. Data-driven surrogates have also been applied to diagnose cooling inefficiency at HPC facilities (Jadhav and Liu, 2026), but these approaches identify waste rather than prescribing constrained optimal control.
@@
Table A1 in Appendix A summarizes the positioning of representative works. No existing paper simultaneously addresses physics-based digital twin modeling, systematic optimization, HPC liquid cooling, explicit practical constraints, and a formal implementability metric.
@@ 1.2 Digital twins for thermal system modeling and optimization
The digital twin concept, originally articulated for product lifecycle management (Grieves and Vickers, 2017) and formalized for industrial applications (Tao et al., 2019), has found growing adoption in data center energy systems. The IISE community has recognized digital twins as a transformative methodology for design and manufacturing (Knapp et al., 2020; Tan and Matta, 2024). For optimization applications where control strategies must be explored beyond the range of historical operating data, physics-based digital twins offer superior generalization and interpretability compared to purely data-driven alternatives (Afram and Janabi-Sharifi, 2014).
@@
The Modelica language and its associated Buildings Library (Wetter et al., 2014) have emerged as the dominant open-source platform for equation-based modeling of HVAC and cooling systems. Modelica-based data center cooling models have been developed for air-cooled facilities (Fu et al., 2019a, b), district cooling systems (Hinkelman et al., 2022), chiller plants with water-side economizers (Fan et al., 2021), and comparative co-design of cooling controls (Grahovac et al., 2023). More recently, the Buildings Library has been coupled with EnergyPlus through the Spawn framework (Wetter et al., 2024).
@@
For HPC cooling specifically, the Exascale Digital Twin (ExaDigiT) project at Oak Ridge National Laboratory developed a digital twin framework for liquid-cooled supercomputers using the TRANSFORM library within the Modelica environment (Brewer et al., 2024; Kumar et al., 2024; Greenwood et al., 2024). However, ExaDigiT focused on verification and validation rather than systematic optimization, a pattern that recurs across the literature: several groups have built physics-based cooling models or digital twin frameworks (Fu et al., 2019a; Hinkelman et al., 2022; Athavale et al., 2024; Sarkar et al., 2024), but the step from validated model to actionable optimization is frequently omitted. The present work constructs an independent Modelica-based digital twin using the Buildings Library (Wetter et al., 2014), purpose-built for steady-state optimization. The model is calibrated against 47,186 validated operational records (Sun et al., 2024) and validated to ASHRAE Guideline 14 (ASHRAE, 2014) criteria, achieving CV-RMSE of 1.96–2.67% and NMBE within $\pm$2.5% across all subloops (Section 2.3).
@@ 1.3 The implementability gap: from theoretical optima to practical control
A persistent challenge in cooling system optimization is the gap between what is theoretically optimal and what can be practically implemented. This gap arises from actuator limitations, thermal inertia, safety margins, and operational protocols requiring gradual system behavior. Despite the practical importance of these constraints, the literature lacks a formal framework for quantifying their effect on achievable energy savings.
Related concepts exist in adjacent domains but do not address cooling system control. The building energy performance gap documents 20–550% discrepancies between design-stage predictions and as-built operation (de Wilde, 2014; Van Dronkelaar et al., 2016), but this concerns modeling fidelity rather than the cost of operational constraints. The simulation-to-real gap in robotics and RL describes controller degradation when transferring from simulation to physical systems (Salvato et al., 2021; Zhao et al., 2020; Chen et al., 2020), but without defining a quantitative metric for cooling applications. The closest quantitative analog comes from energy storage optimization, where ramp-rate constraints reduce arbitrage profit by 10–35%, with constrained solutions retaining 65–90% of unconstrained profit (Hashmi et al., 2024).
We address this gap by introducing the implementability gap, defined as the fraction of theoretical optimal energy savings lost due to practical constraints, and its complement the recovery ratio (Section 2.4.4). The framework is operationalized through a layered optimization: Strategy A optimizes only pump flow to establish a conservative baseline; Strategy B co-optimizes pump flow and supply temperature without operational constraints, revealing the theoretical maximum; and Strategy C re-introduces ramp-rate limits on all actuators, yielding the practically implementable savings. The difference between Strategies B and C directly measures the implementability gap.
@@ 1.4 Contributions and paper organization
    A validated Modelica-based digital twin of the Frontier supercomputer’s three-subloop liquid cooling system, calibrated against one year of operational data (47,186 records) with CV-RMSE of 1.96–2.67%.
@@
    A layered optimization framework (flow-only, unconstrained co-optimization, and ramp-constrained co-optimization) using SLSQP, representing the first application of classical gradient-based optimization to HPC cooling systems.
@@
    The implementability gap metric and recovery ratio, formal measures of the energy savings lost and retained when ramp-rate constraints are imposed on theoretically optimal strategies.
@@
    Operational insights including 20.4% savings from flow-only optimization, 30.1% from unconstrained co-optimization, and 27.8% from ramp-constrained co-optimization (92.4% recovery ratio), demonstrating that the majority of theoretical savings are practically recoverable.
Section 2 presents the digital twin model and optimization formulation. Section 3 presents validation and optimization results. Section 4 discusses insights, robustness, and limitations. Section 5 concludes.
@@ 2.2.1 Heat exchanger model
where $f_{i}$ is the fraction of total flow through subloop ii. This expression follows directly from the first law of thermodynamics applied to the coolant stream passing through the CDU cluster, under the assumption that all waste heat $Q_{i}$ is absorbed by the HTW fluid. The flow fractions are determined by the parallel hydraulic network topology and remain approximately constant at $f_{1}=0.244$, $f_{2}=0.258$, and $f_{3}=0.498$ across the full operating range, as derived from the operational dataset and confirmed by the Modelica hydraulic model (Section 3).
@@
To enable closed-form optimization (Section 2.4), the heat exchanger thermal resistance is characterized by an aggregate overall heat transfer coefficient-area product UAUA. Under the assumption that the HTW-side convective resistance dominates the overall thermal resistance (i.e., wall conduction and CTW-side resistance are comparatively small), the Dittus-Boelter correlation for turbulent internal flow yields $h\propto\dot{m}^{0.8}$ (Incropera et al., 2007), and consequently: